OpenStudy (anonymous):

4x-2(-2x-3)=10x a)-8 b)3 c) no solution d) all real numbers

OpenStudy (anonymous):

b)3

OpenStudy (anonymous):

4x+4x+6=10x 8x+6=10x -8x 6=2x /2 x=3 B is the answer

OpenStudy (anonymous):

4x-2(-2x-3)=10x 4x+4x+6=10x 8x+6=10x | -8x 6 = 2x | :2 3 = x